Calculate the mass of the solute 
 
100cm³ of N/10 KOH

Grade:12

1 Answers

BIKRAM
3 years ago
Here, Solute is KOH and Solvent is water.
Concentration of KOH solution =(N/10) =0.1(N)
Molecular weight of KOH= 39+16+1= 56g/mol
Equivlent weight of KOH= 56g/mol
Amount of solute(KOH) in 100cc solution=(0.1mol/lit x100cc)/1000cc= 0.01mol=(0.01x56)g=0.56g